### v1

@@ 3.4 Theory for read-out persistent currents in cold atoms systems
The persistent current of ultracold atoms is intrinsically related to the particles’ momenta in the condensate. Monitoring the persistent current typically involves measurement of the momentum distribution or interference that highlight the momentum occupation of the particles. From the experimental point of view, this can be done in different ways. These include a non-destructive measurement scheme for the winding number and chirality of the persistent current through cavity optomechanical detection [309, 310], a protocol of coupling the ring to a rectilinear guide [311] and a scheme to measure rotation in a ring by monitoring the precession of the collective acoustic modes [312, 313]. Here we focus on the most common techniques implemented experimentally in the past years, time-of-flight (TOF) and interference dynamics after expansion.
@@
The theoretical approach to TOF accounts to compute the momentum distribution of the system at the instant of time in which the trap is open $t=0$ [318]. The connection between the initial momentum distribution of the particles in the trap and the long-time TOF momentum distribution in position space can be understood in analogy of the Fraunhofer diffraction of light. Here, we sketch such a connection in the lattice (see [319] for the continuous). The one-body correlator is defined as $n(\textbf{r},\textbf{r}^{\prime},t)\equiv\langle\Psi^{\dagger}(\textbf{r},t)\Psi(\textbf{r}^{\prime},t)\rangle$ with r being the coordinate space, and $\Psi^{\dagger}(\textbf{r})$ and $\Psi(\textbf{r})$ are the creation and annihilation operators.
@@
where $\eta_{j}$ provides the width and $\omega_{0}=\frac{\hbar}{m}$ with $\hbar$ and $m$ denoting Planck’s constant and the particles’ mass respectively. Moreover, we note that, within the lattice description, the final momentum distribution is mainly affected by the correlations the particles had before their release from the trap, and the spatial evolution of the falling atoms is encoded on the expansion of the Wannier functions.
For a condensate flowing along ring-shaped circuits, TOF displays a characteristic shape in which the density around k is suppressed. The TOF image (taken from the top of the expanding condensate, along the falling direction) shows a donut shape which is very different from a bell-shaped image in absence of circulation [322], see Fig.9. The value of the donut radius results to change in discrete steps corresponding to the quantization of angular momentum of the condensate [50, 323, 46, 324]. In the non-interacting case, the momentum distribution can be calculated exactly, giving rise to a simple expression that can allow us to understand the existence of the maximum and hole of the density distribution observed in experiments. In particular, $n(\textbf{k})$ is given by [164, 87]:
@@
Figure 9: Momentum distribution for non-interacting particles for (a) zero (b) non-zero current as per (82). (c) Generalized momentum distribution for SU($N$) fermions. For bosonic systems, all particles occupy the same momenta and therefore, the system goes from a maximum to a hole as soon as the system acquires angular momentum. Figures taken from [87].
@@
The interference between the two clouds can be again understood in terms of the mean-field order parameter. Considering that the two clouds are phase coherent, they will interfere depending on their relative phase, $\delta\varphi$, giving a density interference of $n=|\mathbf{\Psi}_{R}+\mathbf{\Psi}_{C}|^{2}=n_{R}+n_{C}+2\sqrt{n_{R,\ell}n_{C}}\cos(\delta\varphi)$, with indices ${R,\,C}$ indicating the ring and center clouds respectively. In the case of interference between two clouds, one wants to avoid the far-field limit as the main contribution will come from the densities and the interference term will be relatively small and thus, hard to detect. To obtain the spiral patterns observed in experiments one needs to consider both the radial and azimuthal parts of the interference. The wavefunction in polar coordinates with radius $r$ and angle $\theta$ is described as $\mathbf{\Psi}_{\ell}(\mathbf{r},t)=\sqrt{n(r,t)}\exp(i\varphi_{\ell}(r,t))\exp(i\ell\theta)$. Here, we have the ring wavefunction, $\mathbf{\Psi}_{R,\ell}$ and the center cloud, $\mathbf{\Psi}_{c}$, which combined gives the interfered density
@@
Fig. 10 displays the interference pattern obtained for the interference between a single particle in the center and one in the ring at given lattice site, as well as the total $G_{R,C}(\mathbf{r},\mathbf{r}^{\prime},t)$ with $\mathbf{r}^{\prime}=(R,0)$. In the same figure, we also show the interference pattern in the form of spirals as well as the dislocation patters obtained when multiple momenta are occupied, obtained through Eq. (87). Similar patterns are visible for interacting system featuring entangled superpositions of momenta [123].
@@ 4.1 Experimental results
Because of difficulties like compensating for drift in the position of the magnetic trap relative to the plug beam, experiments switched to using all-optical ring trap techniques that had been under conceptual and technical development for some time [350, 351, 352, 145, 344, 353, 354], and that had been leading to successful implementations employing “painted” potentials [58] and spatial light modulators [345]. The laser configuration most commonly used to generate an optical ring trap includes a horizontally propagating laser field that mainly provides the vertical confinement, and a patterned vertically propagating beam that provides radial confinement leading to a ring potential or other desired configuration. The experiment at NIST [350] was overhauled to implement an all red-detuned ring trap, with a single horizontal “sheet” beam, and a vertically propagating Laguerre-Gauss beam providing radial confinement. The trap was sufficiently smooth and stable that persistent currents were observed to survive for up to 40 seconds, appearing to be limited only by losses due to collisions with background gas molecules [49]. In these experiments, the persistent current was created using a different two-photon technique than the Bragg method used earlier. Co-propagating beams of differing OAM were used to drive a Raman transition between different magnetic sublevels, instead of using an intermediate state of large linear momentum. This approach required only a single pulse-pair, and resulted in less spurious excitation of the condensate by beam mode imperfections. When a superfluid “weak link” created by a repulsive blue-detuned barrier beam was introduced, the current was observed to decay when the velocity through the barrier approached the estimated local sound speed.
@@
Another important experimental development was the introduction of movable barrier potentials that could be used to drive currents around a ring and/or probe the circulation state of the superfluid in the ring. As described in Sec. 2.3, if one transforms to a rotating frame where the barrier potential is stationary, there is an effective gauge potential and a non-zero “circulation flux” threading through the ring. It was shown to be possible to deterministically and reversibly change the circulation state by driving phase slips with a moving barrier [46, 361], and that the response of the system to a cyclic change of the rotation speed became increasingly hysteretic as the critical current of the moving weak link was increased [360] - see also Sec. 4.4. The hysteresis experiment was conducted with a ring trap employing a blue detuned “dark-ring” potential, instead of a red-detuned “bright-ring”, which was found to have some advantages for creating a more azimuthally uniform ring in the presence of optical imperfections –see Fig. 12. Continuing experiments at the University of Arizona investigated the formation and evolution of turbulent superfluid states involving large numbers of vortices ($\ell>20$) generated by stirring of a wide, flat annular BEC. They found that the system evolved dissipatively toward a persistent current state with some of the vortices being pinned by the central potential, and others leaving the system, and compared their experimental data against numerical simulations showing evidence of an inverse energy cascade consistent with expectations for 2D quantum turbulence [362]
@@
The notable success realizing stable high-order currents in the experiments at Arizona was in part due to the smoothness of their optically plugged magnetic trap. One substantial advantage of time-averaged and RF dressed magnetic traps in general [341, 342, 343, 344, 58, 346, 347] is that they are intrinsically much smoother than optical traps. There has been a string of remarkable recent results leveraging this advantage in experiments using magnetic and and hybrid magnetic/optical traps [373] capable of realizing unusual geometries like rings and bubbles with sizes ranging up to the millimeter scale. One breakthrough experiment conducted at FORTH in Heraklion was able to accelerate a BEC up to supersonic speeds around such a ring using a “bang-bang” scheme, with no evidence of heating as the it continued around the ring for 15 seconds over a total travel distance of 40 cm [80]. Concurrently, experiments at LPL in Paris were carried out where angular momentum selective evaporation of a rotating cloud was used to spin it up to supersonic speeds where it formed a narrow dynamically stabilized annulus 50 $\mu$m in diameter and continued to rotate for more than one minute [374]. The potential of such smooth rf-dressed potentials and waveguides for guided matter-wave interferometer is an important area of continuing investigation.
@@ 4.2 Theoretical understanding - persistent currents in one-dimensional rings
In the case of non-interacting bosons, i.e. upon setting $g=0$ all particles form a mesoscopic Bose-Einstein condensate, and they all occupy the same single-particle energy level. In this case, the total energy is given by
@@
Figure 15: (a) Average angular momentum per particle (main) and ground-state energy (inset) for attractively interacting bosons on a lattice ring as a function of artificial gauge field, from numerical exact diagonalization calculations at varying particle number, for chosen values of interaction strength as indicated on the figure. (b) Ground-state energy (top) and persistent current (bottom) for strongly repulsive SU(2) bosons on a lattice ring as a function of artificial gauge field, from Bethe Ansatz for 4 particles. Figures from [120, 91].
@@
The addition of a localized barrier potential on a ring breaks rotational symmetry and allows to mix angular momentum states. This leads to opening of gaps in the energy landscape in correspondence to the crossing points. The shape of the angular momentum steps gets smeared, and the persistent current amplitude decreases as compared to the clean ring case. In fact, the persistent current amplitude displays a non-monotonous behaviour as a function of the interaction strength, for any value of the barrier strength [121]. This is due to a competition of different effects. At weak interactions, mean-field interactions screen the barrier, which results effectively smaller than its nominal value. As a result, the persistent amplitude increases with interaction strength. Mean-field description breaks down at intermediate interactions, where quantum fluctuations start to play an important role. In this regime, the system can be suitably described by the Luttinger liquid theory. The fluid is described at low-energy by a harmonic-fluid model, recalling Eq. (40)
@@ 4.3 Persistent currents in weakly interacting three-dimensional ring traps
While the one-dimensional analysis can capture relevant traits of the physics of the persistent currents, important features of the phenomenon are distinctive of higher dimensions - see Sec. 4.1. In this case, the theoretical analysis can be conducted through mean-field methods. In particular, the time-dependent Gross-Pitaevskii equation (TDGPE, Eq. (30)) has turned out to provide an excellent approximation to study the system in weak interaction regimes. Typically, the initial condition for the TDGPE is the equilibrium condition in the initial trap and is obtained by first solving the time-independent Gross-Pitaevskii equation, e.g. by evolution in imaginary times. Then, a specific dynamical protocol is implemented and the condensate density $\rho(\mathbf{r},t)$ and phase $\phi((\mathbf{r},t))$ are obtained from the condensate wavefunction by setting $\mathbf{\Psi}(\mathbf{r},t)=\sqrt{\rho(\mathbf{r},t)}e^{i\phi(\mathbf{r},t)}$. Similarly, the momentum distribution is obtained by $n(\mathbf{k},t)=|\mathbf{\tilde{\Psi}}(\mathbf{k},t)|^{2}$, where $\mathbf{\tilde{\Psi}}(\mathbf{k},t)$ is the spatial Fourier transform of the condensate wavefunction.
@@
There are indeed some cases where the GPE does not provide a sufficiently accurate description of the experiment. This is the case when fluctuations, typically of thermal origin, enter into play in the experiment and need to be included in the simulations. In this regime, a useful approach is provided by the Truncated Wigner Approximation as implemented e.g. in Refs. [367, 192]. In essence, this classical field approach includes to lowest order thermal and quantum fluctuations beyond the Gross-Pitaevskii equation. Starting from a 3D Bose-Hubbard Hamiltonian, the method performs a classical field approximation on each site. The fields are propagated by classical equations of motion, with initial conditions generated by sampling from the Wigner distribution of a weakly interacting Bose gas in the Bogoliubov approximation at finite temperature. The resulting classical trajectories are finally averaged over the initial conditions for the calculation of observables. This method has been used to perform theoretical investigations of the decay of supercurrents [367] and to design a protocol to realize an atomtronic SQUID [192].

### v3

@@ 3.4 Theory for read-out persistent currents in cold atoms systems
The persistent current of ultracold atoms is intrinsically related to the particles’ momenta in the condensate. Monitoring the persistent current typically involves measurement of the momentum distribution or interference that highlights the momentum occupation of the particles. From the experimental point of view, this can be done in different ways.
Recent studies have introduced a minimally destructive method to detect persistent currents in a ring-shaped Bose-Einstein condensate (BEC) by coupling it to an optical cavity driven by light carrying orbital angular momentum. This setup allows for real-time, in-situ measurement of the condensate’s rotation by analyzing the cavity transmission spectra, which are sensitive to both the magnitude and sign of the winding number. The technique has been demonstrated to detect not only persistent currents but also soliton dynamics within the condensate [333, 334, 335]. Additionally, an approach involving coupling a ring-shaped condensate to a straight waveguide was also considered as a way to investigate currents in these systems. By observing the dynamics of atoms transitioning between the ring and the guide, one can infer the winding number of the condensate through the specific population transfer in the guide [336]. Further techniques monitor the precession or Doppler shift of collective acoustic modes in the BEC. Preparing excited counter-propagating sound waves, under the presence of a rotating condensate, leads to a precession of the density modulation that can be monitored by in-situ measurements [337, 338]. Here we focus on the most common techniques implemented experimentally in the past years, time-of-flight (TOF) and interference dynamics after expansion.
@@
The theoretical approach to TOF accounts to compute the momentum distribution of the system at the instant of time in which the trap is open $t=0$ [344]. The connection between the initial momentum distribution of the particles in the trap and the long-time TOF momentum distribution in position space can be understood in analogy of the Fraunhofer diffraction of light. Here, we sketch such a connection in the lattice (see [345] for the continuous case). The one-body correlator is defined as $n(\textbf{r},\textbf{r}^{\prime},t)\equiv\langle\Psi^{\dagger}(\textbf{r},t)\Psi(\textbf{r}^{\prime},t)\rangle$ with r being the coordinate space, and $\Psi^{\dagger}(\textbf{r})$ and $\Psi(\textbf{r})$ are the creation and annihilation operators.
@@
where $\eta_{j}$ provides the width and $\omega_{0}=\frac{\hbar}{m}$ with $\hbar$ and $m$ denoting Planck’s constant and the particles’ mass respectively. Moreover, we note that, within the lattice description, the final momentum distribution is mainly affected by the correlations the particles had before their release from the trap, and the spatial evolution of the falling atoms is encoded on the expansion of the Wannier functions. For a condensate flowing along ring-shaped circuits, TOF displays a characteristic shape in which the density around k is suppressed. The TOF image (taken from the top of the expanding condensate, along the falling direction) shows a donut shape which is very different from a bell-shaped image in absence of circulation [187], see Fig.9. The value of the donut radius results to change in discrete steps corresponding to the quantization of angular momentum of the condensate [54, 348, 50, 349]. In the non-interacting case, the momentum distribution can be calculated exactly, giving rise to a simple expression that can allow us to understand the existence of the maximum and hole of the density distribution observed in experiments. In particular, $n(\textbf{k})$ is given by [177, 90]:
@@
(a)(b)(c)
Figure 9: Momentum distribution (top) and its cross-section (bottom) for non-interacting particles for (a) zero and (b)/(c) non-zero angular momentum as per Eq. (76). For bosonic systems, all particles occupy the same momenta and therefore, the system goes from a maximum to a hole as soon as the system acquires angular momentum. However, for fermions this is only the case when the number of particles is equal to the number of components as will be discussed in Sec. 5. Figures taken from [90].
@@
The interference between the two clouds can be again understood in terms of the mean-field order parameter. Considering that the two clouds are phase coherent, they will interfere depending on their relative phase, $\delta\varphi$, giving a density interference of $n=|\mathbf{\Psi}_{R}+\mathbf{\Psi}_{C}|^{2}=n_{R}+n_{C}+2\sqrt{n_{R,\ell}n_{C}}\cos(\delta\varphi)$, with indices ${R,\,C}$ indicating the ring and center clouds respectively, where $\ell$ corresponds to the angular momentum. In the case of interference between two clouds, one wants to avoid the far-field limit as the main contribution will come from the densities and the interference term will be relatively small and thus, hard to detect. To obtain the spiral patterns observed in experiments one needs to consider both the radial and azimuthal parts of the interference. The wavefunction in polar coordinates with radius $r$ and angle $\theta$ is described as $\mathbf{\Psi}_{\ell}(\mathbf{r},t)=\sqrt{n(r,t)}\exp(i\varphi_{\ell}(r,t))\exp(i\ell\theta)$. Here, we have the ring wavefunction, $\mathbf{\Psi}_{R,\ell}$ and the center cloud, $\mathbf{\Psi}_{c}$, which combined gives the interfered density
@@
Fig. 10 displays the interference pattern obtained for the interference between a single particle in the center and one in the ring at given lattice site, as well as the total $G_{R,C}(\mathbf{r},\mathbf{r}^{\prime},t)$ with $\mathbf{r}^{\prime}=(R,0)$. In the same figure, we also show the interference pattern in the form of spirals as well as the dislocation patterns obtained when multiple momenta are occupied, obtained through Eq. (81). Similar patterns are visible for interacting system featuring entangled superpositions of momenta [129].
@@ 4.1 Experimental results
Because of difficulties like compensating for drift in the position of the magnetic trap relative to the plug beam, experiments switched to using all-optical ring trap techniques that had been under conceptual and technical development for some time [374, 375, 376, 151, 368, 377, 378], and that had been leading to successful implementations employing “painted” potentials [62] and spatial light modulators [369]. The laser configuration most commonly used to generate an optical ring trap includes a horizontally propagating laser field that mainly provides the vertical confinement, and a patterned vertically propagating beam that provides radial confinement leading to a ring potential or other desired configuration. The experiment at NIST [374] was overhauled to implement an all red-detuned ring trap, with a single horizontal “sheet” beam, and a vertically propagating Laguerre-Gauss beam providing radial confinement. The trap was sufficiently smooth and stable that persistent currents were observed to survive for up to 40 seconds, appearing to be limited only by losses due to collisions with background gas molecules [53]. In these experiments, the persistent current was created using a different two-photon technique than the Bragg method used earlier. Co-propagating beams of differing OAM were used to drive a Raman transition between different magnetic sublevels, instead of using an intermediate state of large linear momentum. This approach required only a single pulse-pair instead of two pairs, and resulted in less spurious excitation of the condensate by beam mode imperfections. When a superfluid “weak link” created by a repulsive blue-detuned barrier beam was introduced, the current was observed to decay when the velocity through the barrier approached the estimated local sound speed.
@@
Another important experimental development was the introduction of movable barrier potentials that could be used to drive currents around a ring and/or probe the circulation state of the superfluid in the ring. As described in Sec. 2.3, if one transforms to a rotating frame where the barrier potential is stationary, there is an effective gauge potential and a non-zero “circulation flux” threading through the ring. It was shown to be possible to deterministically and reversibly change the circulation state by driving phase slips with a moving barrier [50, 385], and that the response of the system to a cyclic change of the rotation speed became increasingly hysteretic as the critical current of the moving weak link was increased [384] - see also Sec. 4.4. The hysteresis experiment was conducted with a ring trap employing a blue detuned “dark-ring” potential, instead of a red-detuned “bright-ring”, which was found to have some advantages for creating a more azimuthally uniform ring in the presence of optical imperfections –see Fig. 12. Continuing experiments at the University of Arizona investigated the formation and evolution of turbulent superfluid states involving large numbers of vortices ($\ell>20$) generated by stirring of a wide, flat annular BEC. They found that the system evolved dissipatively toward a persistent current state with some of the vortices being pinned by the central potential, and others leaving the system, and compared their experimental data against numerical simulations showing evidence of an inverse energy cascade consistent with expectations for 2D quantum turbulence [386].
@@
The notable success realizing stable high-order currents in the experiments at Arizona was in part due to the smoothness of their optically plugged magnetic trap. One substantial advantage of time-averaged and RF dressed magnetic traps in general [365, 366, 367, 368, 62, 370, 371] is that they are intrinsically much smoother than optical traps. There has been a string of remarkable recent results leveraging this advantage in experiments using magnetic and hybrid magnetic/optical traps [397] capable of realizing unusual geometries like rings and bubbles [398] with sizes ranging up to the millimeter scale. One breakthrough experiment conducted at FORTH in Heraklion was able to accelerate a BEC up to supersonic speeds (16$\times$ their speed of sound) around such a ring using a “bang-bang” scheme(10)(10)(10)A “bang-bang” scheme is a protocol derived from optimal control, characterized by abrupt discrete switching between two states or extreme values [399]. A common example of such a scheme is the discrete Heaviside function. In this particular experiment, the system experiences rapid pulses that act as strong discrete kicks to attain the desired state., with no evidence of heating as it continued around the ring for 15 seconds over a total travel distance of 40 cm [83]. Concurrently, experiments at LPL in Paris were carried out where angular momentum selective evaporation of a rotating cloud was used to spin it up to supersonic speeds where it formed a narrow dynamically stabilized annulus 50 $\mu$m in diameter and continued to rotate for more than one minute as shown in [400]. The potential of such smooth rf-dressed potentials and waveguides for guided matter-wave interferometer is an important area of continuing investigation.
@@ 4.2 Theoretical understanding - persistent currents in one-dimensional rings
In the case of non-interacting bosons, i.e. upon setting $g=0$ all particles form a Bose-Einstein condensate, and they all occupy the same single-particle energy level. In this case, the total energy is given by
@@
Figure 15: (a) Average angular momentum per particle (main) and ground-state energy $E$ (inset) for bosons with attractive interactions $U$ on a lattice ring as a function of artificial gauge field $\Omega$, from numerical exact diagonalization calculations at varying particle number $N$, for chosen values of interaction strength as indicated in the figure. (b) Ground-state energy (top) and persistent current (bottom) for SU(2) bosons on a lattice ring with strong repulsive interactions $c$ as a function of the dimensionless synthetic gauge field $\tilde{\Phi}$, from Bethe Ansatz for 4 particles. The angular momentum $\ell$ and the Young tableau reflecting the symmetry of the ground-state for each piece-wise parabola are shown for bosons (blue) and fermions (orange). Figures from [126, 94].
The addition of a localized barrier potential on a ring breaks rotational symmetry and allows to mix angular momentum states. This leads to opening of gaps in the energy landscape in correspondence to the crossing points. The shape of the angular momentum steps gets smeared, and the persistent current amplitude decreases as compared to the clean ring case. In fact, the persistent current amplitude displays a non-monotonous behaviour as a function of the interaction strength, for any value of the barrier strength [127]. This is due to a competition of different effects. At weak interactions, mean-field interactions screen the barrier, which results in its strength effectively smaller than its nominal value. As a result, the persistent amplitude increases with interaction strength. Mean-field description breaks down at intermediate interactions, where quantum fluctuations start to play an important role. In this regime, the system can be suitably described by the Luttinger liquid theory. The fluid is described at low-energy by a harmonic-fluid model, recalling Eq.(36)
@@ 4.3 Persistent currents in weakly interacting three-dimensional ring traps
While the one-dimensional analysis can capture relevant traits of the physics of the persistent currents, important features of the phenomenon are distinctive of higher dimensions - see Sec. 4.1. In this case, the theoretical analysis can be conducted through mean-field methods. In particular, the time-dependent Gross-Pitaevskii equation (TDGPE, Eq. (27)) has turned out to provide an excellent approximation to study the system in weak interaction regimes. Typically, the initial condition for the TDGPE is the equilibrium condition in the initial trap and is obtained by first solving the time-independent Gross-Pitaevskii equation, e.g. by evolution in imaginary times. Then, a specific dynamical protocol is implemented and the condensate density $\rho(\mathbf{r},t)$ and phase $\phi(\mathbf{r},t)$ are obtained from the condensate wavefunction by setting $\Psi(\mathbf{r},t)=\sqrt{\rho(\mathbf{r},t)}e^{i\phi(\mathbf{r},t)}$. Similarly, the momentum distribution is obtained by $n(\mathbf{k},t)=|{\tilde{\Psi}}(\mathbf{k},t)|^{2}$, where ${\tilde{\Psi}}(\mathbf{k},t)$ is the spatial Fourier transform of the condensate wavefunction.
@@
There are indeed some cases where the GPE does not provide a sufficiently accurate description of the experiment. This is the case when fluctuations, typically of thermal origin, enter into play in the experiment and need to be included in the simulations. In this regime, a useful approach is provided by the Truncated Wigner Approximation as implemented e.g. in Refs. [391, 212]. In essence, this classical field approach includes to lowest order thermal and quantum fluctuations beyond the Gross-Pitaevskii equation. Starting from a 3D Bose-Hubbard Hamiltonian, the method performs a classical field approximation on each site. The fields are propagated by classical equations of motion, with initial conditions generated by sampling from the Wigner distribution of a weakly interacting Bose gas in the Bogoliubov approximation at finite temperature. The resulting classical trajectories are finally averaged over the initial conditions for the calculation of observables. This method has been used to perform theoretical investigations of the decay of supercurrents [391] and to design a protocol to realize an atomtronic SQUID [212]. Other approaches investigating both formation and decay of persistent currents, including stochastic projected GPE, stochastic GPE, stochastic Landau-Ginzburg, Bogoliubov and others [405, 350, 360, 362, 379, 391, 387, 392, 390, 116, 393, 394, 406].